OpenStudy (anonymous):

john is four years older than frank, the sum of their ages is 36 How would this look as an equation?

Directrix (directrix):

Let J represent Frank's age and F represent Frank's age. So, J = 4 + F. >the sum of their ages is 36 J + F = 36 Substituting, 4 + F + F = 36. --> Solve for F @mel12561

Directrix (directrix):

Let me know what you get and we'll compare answers, okay?